What is the distribution of electrons in an oxygen atom?

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Jayesh Vashishtha
2 years ago
Dear student,
 
The total number of electrons present in oxygen is 8.
Therefore, for O electron configuration will be 1s22s22p4.
The distribution of electrons in oxygen atom is given by:

First orbit or K shell =2(2n^2=2×1^1=2)
Second orbit or L shell =6(8−2=6)
 
regards,
Team Askiitians